OpenStudy (nightmarenight):

The endpoints of a line segment are at the coordinates (3, 4, 6) and (5, 7, 2). What is the midpoint of the segment?

OpenStudy (bibby):

so basically \((\dfrac{3+5}{2},\dfrac{4+7}{2},\dfrac{6+2}{2})\)

OpenStudy (nightmarenight):

8/2 , 11/2 , 8/2???

OpenStudy (bibby):

aye aye

OpenStudy (bibby):

4,5.5,4

OpenStudy (nightmarenight):

these are my options (4, 1.5, 2) (7, 18, -10) (1, -15, 14) (1.5, -2, 3)

OpenStudy (bibby):

that doesn't make much sense to me, hold on

OpenStudy (nightmarenight):

Alrighty

OpenStudy (bibby):

are you sure the question is for those choices

OpenStudy (bibby):

o rofl there are negatives you silly billy

OpenStudy (nightmarenight):

Huh? OH in the question??? D: i didn even see

OpenStudy (bibby):

\((\dfrac{3+5}{2},\dfrac{-4+7}{2},\dfrac{6+-2}{2})\)

OpenStudy (nightmarenight):

8, 3, 4

OpenStudy (nightmarenight):

wait divided by 2???

OpenStudy (nightmarenight):

Oh It's a

OpenStudy (bibby):

yeah

OpenStudy (bibby):

good job son i am proud